OpenStudy (anonymous):

If K = {(x, y )|x - y = 5}, find the corresponding range of y for the domain {0, 2, 4}. {-5, -3, -1} {0, 2, 4} {5, 7, 9}

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

Domain =x Range = y

OpenStudy (anonymous):

Yes, I know this. Lol

OpenStudy (anonymous):

So for each value of x what would be the y value?

OpenStudy (anonymous):

That's what I don't understand

OpenStudy (anonymous):

put x=0,2,4 in x-y=5,find y for each value

OpenStudy (anonymous):

I don't understand the x-y=5 part and how I'm supposed to find the value of y from that

OpenStudy (anonymous):

x-y=5 if x=0 0-y=5 y=-5 similarly find other values.

OpenStudy (anonymous):

#1

OpenStudy (anonymous):

if x=2 2-y=5 -y=5-2 -y=3 y=-3

OpenStudy (anonymous):

if x=4 y=?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

4-y=5 -y=5-4 -y=1 y=-1

OpenStudy (anonymous):

If K = {(x, y ) | x - y = 5}, is Set K a function? Yes No

OpenStudy (anonymous):

yes,because for one value of x we get only one value of y.